Currently, Warren Industries can sell 15 – year, $1,000-par-value bonds paying annual interest at a 9% coupon rate. Because current market rates for similar bonds are just under 9%, Warren can sell its bonds for $1,040 each; Warren will incur flotation costs of $25 per bond. The firm is the 21% tax bracket.

a. The net proceeds from the sale of the​ bond, Upper N Subscript d is$               . ​(Round to the nearest​ dollar.)

b. Using the​ bond's YTM, the​ before-tax cost of debt is                  %. ​(Round to two decimal​ places.)

Using the​ bond's YTM, the​ after-tax cost of debt is                       %.​(Round to two decimal​ places.)

c. Using the approximation​ formula, the b​efore-tax cost of debt is            %. (Round to two decimal​ places.)

Using the approximation​ formula, the after​-tax cost of debt is                  %.​(Round to two decimal​ places.)
Answer #1

a)
Net Proceed = Current Price - Flotation Cost
= $1,040 - $25
= $1,015.

Net proceed = $1,015.

b)
FV = 1000
Nper = 15
PMT = 1000 * 9% = 90
PV = 1015

Before tax cost of debt can be calculated by using the following excel formula:
=RATE(nper,pmt,pv,fv)
=RATE(15,90,-1015,1000)
= 8.82%

Before tax cost of debt = 8.82%

After tax cost of debt = Before tax cost of debt * (1 - tax rate)
= 8.82% * (1 - 0.21)
= 6.96%

After tax cost of debt = 6.96%

c)
Approximate YTM = [C + (F - P) / n] / [(F + P) / 2]
= (90 + (1000 - 1015) / 15) / ((1000 + 1015) / 2)
= 89 / 1007.5
= 8.83%

Before tax cost of debt = 8.83%

After tax cost of debt = 8.83% * (1 - 0.21) = 6.98%

After tax cost of debt = 6.98%